# Surg-3M: A Dataset and Foundation Model for Perception in Surgical Settings

### v2

@@ Appendix D Downstream Tasks: Data Splits and Fine-tuning
Fine-tuning on Downstream tasks. We used the teacher model from SurgFM as our backbone and employed three different protocols for downstream task fine-tuning. The images were resized to $448\times 448$ for all downstream tasks. The downstream models with linear heads were trained with a batch size of 112, an initial learning rate (LR) of 1e-4, the AdamW optimizer, and cross-entropy loss. An early stopping criterion was applied, stopping training after 10 epochs if the validation loss shows no improvement.
@@
For the surgical tool presence detection tasks and the CholecT50 action recognition task, we used a linear head. For the surgical phase recognition tasks, we utilized a TCN [12] head and followed the two-stage training approach outlined in TeCNO [12]. In the first stage, we trained the backbone with a linear head to perform frame-wise classification without temporal context. Then, we employed the TCN head to incorporate temporal information of the extracted features for predictions. The default TCN configuration was used, consisting of two stages, each containing nine layers. The TCN head was trained using the Adam optimizer with an initial learning rate of 5e-3 for 70 epochs to ensure sufficient training. The checkpoint with the lowest validation loss was selected as the optimal checkpoint. For the SAR-RARP50 [48] action recognition task, we trained our model in two modes: online, where the SurgFM teacher model was used as backbone with a TCN head that could see only the information from previous frames; and offline, where the SurgFM teacher model was paired with an MS-TCN head [33] to incorporate both previous and future frame information for a fair comparison with the methods in [48]. For the TCN head, we maintained the same model configuration and hyperparameters as in the surgical phase recognition tasks. For the MS-TCN head, we adopted a 3-stage, 11-layer configuration, following the basic model setup in [33]. The model was trained using the Adam optimizer with an initial learning rate of 5e-4 for 70 epochs.
Data scaling. In Sec. 6 of the manuscript, we present the results of the data scaling experiments. For each scale (10%, 25%, 50%), we conducted a 5-fold cross-validation, with each fold randomly selected from the training dataset. The model configurations and hyperparameters were kept consistent across different data scales using the same settings as those used for the full training dataset.
Appendix E Evaluation of SurgFM-Vid
@@
To evaluate the performance of our video classifier, SurgFM-Vid, we have split the Surg-3M dataset by videos, with 3369 videos for training and 825 videos for testing. Our vanilla SurgFM was trained on all the Surg-3M frames, therefore, to evaluate the video classification performance fairly, we have trained a new SurgFM model from scratch using only the frames of the 3369 videos in the training set that are accessible to all other classifiers in the comparison (i.e., 2.6 million frames instead of the usual 3.3 million frames in Surg-3M).
@@ Appendix F Dataset and Licensing
We adhere to the same practices as other datasets created from YouTube and various sources on the Web, such as ImageNet [13], Kinetics [9], YouTube-VIS [61], YouTube-8M [1], Insect-1M [41], Moments-in-Time [40], Tai-Chi-HD [51], HD-Villa-100M [60], and AVSpeech [15]. Specifically, we provide the code to download and generate the Surg-3M dataset, a list of links to the original YouTube videos, and the corresponding annotations. Researchers working in academic institutions can request direct access to the full Surg-3M dataset in LMDB format for non-commercial purposes. We will also provide an online form for the YouTube video authors to opt out of our Surg-3M dataset. The Surg-3M dataset is provided under the Creative Commons Attribution 4.0 International (CC BY 4.0) license.

### v3

@@ D.3 Full fine-tuning
We used the teacher model from LemonFM as our backbone and employed three different protocols for downstream task fine-tuning. The images were resized to $\times 224$ 224 for all downstream tasks. The downstream models with linear heads were trained with a batch size of 112, an initial learning rate (LR) of 1e-4, the AdamW optimizer, random seed of 30, and cross-entropy loss. We use fivefold cross-validation for 50% shot fine-tuning. An early stopping criterion was applied, stopping training after 10 epochs if the validation loss shows no improvement.
Surgical phase recognition To compare with other specialist models that are specifically tailored for this task, we adapt the video-level accuracy and Jaccard in the full fine-tuning setting. The accuracy for a video is computed by dividing the number of frames whose class has been correctly predicted by the total number of frames. The overall accuracy for a dataset is defined as the mean accuracy value over all videos, henceforth referred to as video-level accuracy [28, 29, 35, 27]. The Jaccard is computed for each class and video, then averaged: first within videos, and secondly across classes [28, 29, 35, 27].
@@
During training, we utilized a TCN [12] head and followed the two-stage training approach outlined in TeCNO [12]. In the first stage, we trained the backbone with a linear head to perform frame-wise classification without temporal context. Then, we employed the TCN head to incorporate temporal information of the extracted features for predictions. The default TCN configuration was used, consisting of two stages, each containing nine layers. The TCN head was trained using the Adam optimizer with an initial learning rate of 5e-3 for 70 epochs to ensure sufficient training. The checkpoint with the lowest validation loss was selected as the optimal checkpoint.
Appendix E Evaluation of LemonFM-Vid
@@
To evaluate the performance of our video classifier, LemonFM-Vid, we have split the LEMON dataset by videos, with 3369 videos for training and 825 videos for testing. Our vanilla LemonFM was trained on all the LEMON frames, therefore, to evaluate the video classification performance fairly, we have trained a new LemonFM model from scratch using only the frames of the 3369 videos in the training set that are accessible to all other classifiers in the comparison.
@@ Appendix F Dataset and Licensing
We adhere to the same practices as other datasets created from YouTube and various sources on the Web, such as ImageNet [13], Kinetics [9], YouTube-VIS [56], YouTube-8M [1], Insect-1M [42], Moments-in-Time [41], Tai-Chi-HD [49], HD-Villa-100M [55], and AVSpeech [14]. Specifically, we provide the code to download and generate the LEMON dataset, a list of links to the original YouTube videos, and the corresponding annotations. Researchers working in academic institutions can request direct access to the LEMON dataset (1 fps version) in LMDB format for non-commercial purposes. We will also provide an online form for the YouTube video authors to opt out of our LEMON dataset. The LEMON dataset is provided under the Creative Commons Attribution 4.0 International (CC BY 4.0) license.